OpenStudy (anonymous):

Please HELP!!!!!!! Using the function y = 5sin3(x – 25) + 4, answer the following questions: A) What is the maximum value for this graph? B) What is the period for this graph? C) What is the phase shift for this graph?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lets take a couple easy steps to find the maximum what is the maximum of sine ?

OpenStudy (anonymous):

maximum value of sin 0 =1

OpenStudy (misty1212):

yes so if the biggest sine can be is one, whats is the biggest \(5\times\sin(\theta)\)?

OpenStudy (anonymous):

1

OpenStudy (misty1212):

no, you are multiplying it by 5

OpenStudy (anonymous):

Oh..then its 5

OpenStudy (misty1212):

right

OpenStudy (misty1212):

and then add 4 what is the biggest \[5\sin(\theta)+4\] can be?

OpenStudy (anonymous):

9

OpenStudy (misty1212):

done

OpenStudy (anonymous):

That was easy..lol. Thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

you still need the period, which is \[\frac{2\pi}{b}\] in your case \(b=3\)

OpenStudy (anonymous):

Yea. I'm not sure about the period

OpenStudy (misty1212):

did you understand my answer ?

OpenStudy (anonymous):

I didn't see it come up. It said math processing error

OpenStudy (misty1212):

refresh your browswer

OpenStudy (anonymous):

Oh ok I see it now!

OpenStudy (anonymous):

Ok so if i calculate that I should find the period right?

OpenStudy (anonymous):

What about the phase shift?

OpenStudy (misty1212):

just write the period, don't "calculate" it \[\frac{2\pi}{b}=\frac{2\pi}{3}\]

OpenStudy (anonymous):

Ok yea I just realized that

OpenStudy (misty1212):

the phase shift is that 25 out at the end

OpenStudy (anonymous):

Oh ok. Makes sense. Thank you soo much for helping me!!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]